The speedometer readings in passenger cars are based on angular speed of the cars tire axle. Because cars tires are filled
Answer #1

Hey!

This is simple question you just have to read the question carefully 2-3 times.

given that at temp T(c) = 2.80 degree actual speed is 66.4 mph but the speedometer speed is 68.8 mph

and speedometer reading is based on the rotation of axle. and given that speedometer is tuned in summer at temp T(h). that means the distance travel in summer in a particular time by the tyre is calculated and according to that speedometer is tuned.

when the speedometer reading is 68.8 mph that means if temp was T(h) then this 68.8 reading is true reading. radius at T(h) is R(h). let at this speed rpm of axle = N rpm

so speed = 2*pi*R(h)*N/60 = 68.8 mph...................................................equation 1

but the actual speed is 66.4 because of temp is T(c) which is lower than the T(h) so now the radius is R(c).

so speed is = 2*pi*R(c)*N/60 = 66.4 mph..............................................equation 2

divide equation 1 by 2

R(h)/R(c) = 68.8/66.4

R(h) = 0.3279 m

now given that tyre is sealed and due to temp change only radius is changed (pressure, width are constant)

according to ideal gass equation

P*V = mRT

is tyre is sealed mass of air will remain constant

P*V(h)/R*T(h) = P*V(c)/R*T(c)

V(h)/T(h) = V(c)/T(c) ........................................................equation 3

V(h) = pi*[R(h) - R(rim)]^2*(width)

V(c) = pi*[R(c) - R(rim)]^2*(width)

putting values in equation 3

[R(h) - R(rim)]^2 / [R(c) - R(rim)]^2 = T(h)/T(c)

R(rim) =0.1905 m

putting all the values

1.1891 = T(h)/T(c)

T(h) = 1.1891*2.80 = 3.3295 degree